Memofax, Inc. produces memory enhancement software for computers. Sales have been very erratic, with some months showing a profit and some months showing a loss. The company’s contribution format income statement for the most recent month is given below:

  

  Sales (22,500 units at $30 per unit) $ 675,000   
  Less: Variable expenses 472,500   
  Contribution margin 202,500   
  Less: Fixed expenses 211,500   
  Net operating loss $ (9,000)
Required:
1.

Compute the company’s CM ratio and its break-even point in both units and dollars.

      

2.

The sales manager feels that an $26,000 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in a $160,000 increase in monthly sales. If the sales manager is right, what will be the effect on the company’s monthly net operating income or loss? (Use the incremental approach in preparing your answer.)

    

3.

Refer to the original data. The president is convinced that a 10% reduction in the selling price, combined with an increase of $75,000 in the monthly advertising budget, will double unit sales. What will the new contribution format income statement look like if these changes are adopted?

    

4.

Refer to the original data. The company’s advertising agency thinks that a new package would help sales. The new package being proposed would increase packaging costs by $0.5 per unit. Assuming no other changes, how many units would have to be sold each month to earn a profit of $9,500? (Do not round intermediate calculations.)

      

  

5.

Refer to the original data. By automating, the company could slash its variable expenses in half. However, fixed costs would increase by $81,000 per month.

a.

Compute the new CM ratio and the new break-even point in both units and dollars. (Do not round intermediate calculations. Round "Contribution Margin Ratio" to 2 decimal places.)

            

b.

Assume that the company expects to sell 27,000 units next month. Prepare two contribution format income statements: one assuming that operations are not automated, and one assuming that they are. (Do not round intermediate calculations. Round "Per Unit" and "Percentage" to 2 decimal places.)
Answer #1

ANSWER

REQUIRED 1

Contribution margin (CM) ratio = Contribution margin / Sales

= 202,500   / 675,000   

= 0.3

------------------------------------------
Variable expenses per unit = Variable expenses / Number of units

= 472,500   / 22,500

= 21

Contribution margin per unit = Selling price per unit - Variable expenses per unit

= 30 - 21

= 9

Breakeven point in units = Fixed expenses / Contribution margin per unit

= 211,500   / 9

= 23,500 units

------------------------------------------

Breakeven point in dollars = Fixed costs / Contribution margin ratio

= 211,500   / 0.3

= 705,000

------------------------------------------------------------------------------------

REQUIRED 2

Increase in sales = 160,000

Increase in expenses = 26,000

Effect on company's monthly net operating income = Increase in sales - Increase in expenses

= 160,000 - 26,000

= 134,000

Operating income increases by 134,000.

------------------------------------------------------------------------------------

REQUIRED 3

Contribution format Income Statement

Sales [45,000units * 27 per unit] 1,215,000
Variable expenses [45,000 units * 21 per unit] 945,000
Contribution margin 472,500   
Fixed expenses [211,500   +75,000 ] 286,500
Net operating loss (148,000)

* Units sold = 22,500 * 2 = 45,000

* Selling price per unit = 30 - (30*10%) = 27

* Variable expenses per unit = 472,500   / 22,500 = 21

------------------------------------------------------------------------------------

REQUIRED 4

Selling price per unit = 30

Variable expenses per unit = 472,500    / 22,500 = 21

Increased variable expenses per unit = 21 + 0.5 = 21.5

Contribution margin per unit = Selling price per unit - Variable expenses per unit

= 30 - 21.5

= 8.5

Units to be sold to earn desired profit = (Fixed expenses + Desired profit) / Contribution margin per unit

= (211,500+ 9,500) / 8.5

= 26,000 units